jigglypuff314 (jigglypuff314):

Water is running out of a conical funnel at the rate of 3 cubic inches per sec. If the funnel is 12 inches tall and has a radius of 4 inches, find the rate at which the water level is dropping when the water in the funnel is 8 inches below the rim of the funnel...

jigglypuff314 (jigglypuff314):

I'm just confused, could some one point me of what to solve for? like dV/dt or dh/dt or idk :/ and maybe what some terms mean? like does the "3 cubic inches per sec" mean that dV/dt = 3 in^3/s ? or no?

jigglypuff314 (jigglypuff314):

@hartnn @SolomonZelman ? if you're not too busy, could you help please?

OpenStudy (solomonzelman):

You tagged me with hartnn as if I am on the same level. idk this staff, sorry.

hartnn (hartnn):

"3 cubic inches per sec" mean that dV/dt = 3 in^3/s" YES

jigglypuff314 (jigglypuff314):

okay, but then what would I solve for?

hartnn (hartnn):

what is the volume of cone formula ?

jigglypuff314 (jigglypuff314):

V = 1/3 pi r^2 h

hartnn (hartnn):

the radius remains constant, just the height of water changes

hartnn (hartnn):

so, differentiate V = 1/3 pr r^2 h w.r.t time t, what do u get ?

jigglypuff314 (jigglypuff314):

I'm confused... how would the radius be constant?

hartnn (hartnn):

is the funnel changing its shape or size ?

jigglypuff314 (jigglypuff314):

I suppose not...

hartnn (hartnn):

so the radius of cone will not change

jigglypuff314 (jigglypuff314):

so dV/dt = 1/3 pi (2r) (dr/dt) (dh/dt) ?

hartnn (hartnn):

only the height of water changes which leads to change in volume

hartnn (hartnn):

but but we said r is constant, right ?

jigglypuff314 (jigglypuff314):

so dV/dt = 1/3 pi r^2 dh/dt ?

hartnn (hartnn):

yes!

hartnn (hartnn):

dV/dt = 3 r=4 find dh/dt

hartnn (hartnn):

'the rate at which the water level is dropping' = dh/dt

jigglypuff314 (jigglypuff314):

so dh/dt = 9/16pi ?

hartnn (hartnn):

i suppose yes, except if i mis-read the question ....

jigglypuff314 (jigglypuff314):

yeah, probably, because my answer key says it should be -27/(16pi)

hartnn (hartnn):

yes, i was wondering why we did not use the fact that it was 12 inches before and now its 8 inches in height

hartnn (hartnn):

so, change in height = dh = 8-12

jigglypuff314 (jigglypuff314):

? so dh/dt = -4 ?

jigglypuff314 (jigglypuff314):

or new height = 4 ?

hartnn (hartnn):

i probably did a minor mistake somewhere, let me call someone.... @jdoe0001 @ranga

OpenStudy (ranga):



OpenStudy (ranga):

Find the radius x (or r) as a function of h using the similar triangles ABC and DEC.

jigglypuff314 (jigglypuff314):

x = 4/3

OpenStudy (ranga):

x/4 = h/12 x = 1/3h V = 1/3(pi)x^2h = 1/3(pi)(1/9h^2)h = pi/27 * h^3 dV/dt = (pi/27) * 3h^2 * dh/dt

OpenStudy (ranga):

dV/dt = (pi/9) * h^2 * dh/dt Substitute the values and solve for dh/dt

jigglypuff314 (jigglypuff314):

okay, I got it! Thank You! :D

OpenStudy (ranga):

you are welcome.